Im helping my brother with this problem and i cant get it so maybe you can.
If tim has a total of 180 marbles some plain and some colored. there are 32 more plain marbles then color. how many marbles does he have of each?

Answer by Paul(988)

Let x represent the color marbels.
x+(x+32)=180
2x+32-32=180-32
x=74
74+32=106
Hence, there were 106 plain marbles are 74 color marbels